The electric field has a constant value of 3.1 x 102 V/m and is directed downward. The field is the same everywhere. The potential at a point P within this region is 148 V. Find the potential at the following points: (a) 9.4 x 10-3 m directly above P, (b) 2.7 x 10-3 m directly below P, (c) 4.0 x 10-3 m directly to the right of P.
Answer #1

a) 9.4 x 10-3 m directly above P

V - Vp = E*d

V = Vp + E*d = 148+9.4*10^-3*3100 = 177.14 V

b) 2.7 x 10-3 m directly below P

Vp - V = E*d

V = Vp-E*d = 148-3100*2.7/1000 = 139.63 V

c) 4.0 x 10-3 m directly to the right of P

At the same level the potential will be same as all these points are equipotential points.